# Parsimonious Gaussian mixture models with piecewise-constant eigenvalue profiles

### v1

@@ A.1 Proposition 1: number of parameters for the MPSA models
The parameters of the MPSA model of types $\gamma$ are $\pi_{1},\mu_{1},\Sigma_{1},\dots,\pi_{C},\mu_{C},\Sigma_{C}$, with the constraints $\pi_{c}\in\mathbb{R}_{\geq 0},\sum_{c=1}^{C}\pi_{c}=1$, $\mu_{c}\in\mathbb{R}^{p}$ and $\Sigma_{c}\in\mathrm{Sym}_{+}(\gamma_{c})$. The number of free parameters for $\pi_{1},\dots,\pi_{C}$ (belonging to the standard simplex) is $C-1$, since the last mixture proportion $\pi_{C}$ is automatically determined from the $C-1$ other ones, via $\pi_{C}=1-\sum_{c=1}^{C-1}\pi_{c}$. Then for each $c\in[\![1,C]\!]$, the parameters $\mu_{c}$ and $\Sigma_{c}$ describe a PSA model of type $\gamma_{c}$. Therefore, the number of free parameters is $\kappa(\gamma_{c})$, as defined in Equation (2). Since neither the means nor the covariance matrices are shared between the mixture components, the number of parameters for $\mu_{1},\Sigma_{1},\dots,\mu_{C},\Sigma_{C}$ is simply the sum $\sum_{c=1}^{C}\kappa(\gamma_{c})$. Eventually, since the mixture proportions are independent from the means and covariance matrices, the total number of parameters is ${\kappa(\gamma)=C-1+\sum_{c=1}^{C}\kappa(\gamma_{c})}$. ∎
@@ A.2 Theorem 2: E-step of the EM
Let us define the cost function $K_{c}$ similarly as in [11] via the following relationship: $\ln(\pi_{c}\,\mathcal{N}(\,x\mid\mu_{c},\Sigma_{c}\,))=-\frac{K_{c}(x)+p\ln 2\pi}{2}$. Re-injecting $K_{c}$ in the posterior probabilities yields
@@ A.3 Theorem 3: M-step of the EM
The solutions to such problems are given in the maximum likelihood estimation theorem of [59]. One has $\hat{\mu}_{c}=\frac{1}{n\hat{\pi}_{c}}\sum_{i=1}^{n}t_{ci}\,x_{i}$ and $\hat{\Sigma}_{c}=\sum_{k=1}^{d_{c}}\hat{\lambda}_{ck}(t_{c},\gamma_{c})\,\hat{\Pi}_{ck}(t_{c},\gamma_{c})$, where $\hat{\lambda}_{ck}$ and $\hat{\Pi}_{ck}$ are the block-averaged eigenvalues and eigenspaces of $S_{c}$. ∎
@@ A.4 Theorem 4: monotonicity of the CPEM
The two previous equations (45, 46) simply mean that a classical EM iteration for the MPSA model of types $\gamma$ increases the expected complete-data penalized log-likelihood $\mathbb{E}_{\,y\mid x,\theta\,}[\ln\mathcal{L}_{\mathrm{c}}(\cdot)]-\alpha\,\kappa(\gamma)$, and that an adapted change of types from $\gamma$ to $\gamma^{\prime}$ increases this quantity even more. Using Gibbs’ inequality as classically done in the proof of the EM monotonicity [27], one gets
@@ A.5 Proposition 5: GMM-based single-image denoising
The Proposition 1 of [36] proves that $\mathbb{E}(\,x\mid\tilde{x}\,)=\sum_{c=1}^{C}w_{c}\psi_{c}(\tilde{x})$ with ${}{\psi_{c}(\tilde{x})\coloneqq\mu_{c}+(I_{p}-\sigma^{2}{\Sigma_{c}}^{-1})(\tilde{x}-\mu_{c})}$. Then, using the eigendecomposition of the sample covariance matrices, $\Sigma_{c}=\sum_{k=1}^{d_{c}}\lambda_{ck}\Pi_{ck}$, where $\lambda_{ck}$ are the (ordered decreasing) eigenvalues of $\Sigma_{c}$ and $\Pi_{ck}$ are the associated eigenspaces, one can rewrite:
$\psi_{c}(\tilde{x})$ $={\mu_{c}+\left(I_{p}-\sigma^{2}\sum_{k=1}^{d_{c}}\lambda_{ck}^{-1}\Pi_{ck}\right)(\tilde{x}-\mu_{c})}$ (49)
@@ Appendix B Experiments
This section provides more details about the experiments. The code is written in Python 3.9 and run on an Intel® Core™ i7-10850H CPU with 16 GB of RAM. The code for the EM algorithm is a modification of Mathieu Fauvel’s code (license: GPL-3.0), itself inspired from the R package of [4] (license: GPL-2). Some parts of the code related to PSA inference and model selection were inspired from the code of [59] (license: MIT). The code for the denoising pipeline is a modification of Antoine Houdard’s code (license: MIT).
@@
The used packages are the following: scikit-learn [53] (license: BSD 3), numpy [38] (license: BSD), scipy [64] (license: BSD 3), matplotlib [39] (license: custom), ucimlrepo (license: MIT), scikit-image [68] (license: BSD 3). The code for sampling from a skew-normal distribution comes from Gregory Gundersen’s blog post (license: not specified).
@@
Some datasets come from the UCI Machine Learning repository: wine [2] (license: CC BY 4.0), breast [67] (license: CC BY 4.0) and ionosphere [60] (license: CC BY 4.0). The Simpson image, which is notably used in [36] comes from the website page of Antoine Houdard (license: not specified).

### v2

@@ A.1 Proposition 1: number of parameters for the MPSA models
The parameters of the MPSA model of types $\gamma$ are $\pi_{1},\mu_{1},\Sigma_{1},\dots,\pi_{C},\mu_{C},\Sigma_{C}$, with the constraints $\pi_{c}\in\mathbb{R}_{\geq 0},\sum_{c=1}^{C}\pi_{c}=1$, $\mu_{c}\in\mathbb{R}^{p}$ and $\Sigma_{c}\in\mathrm{Sym}_{+}(\gamma_{c})$. The number of free parameters for $\pi_{1},\dots,\pi_{C}$ (belonging to the standard simplex) is C1C-1, since the last mixture proportion $\pi_{C}$ is automatically determined from the C1C-1 other ones, via $\pi_{C}=1-\sum_{c=1}^{C-1}\pi_{c}$. Then for each $c\in[\![1,C]\!]$, the parameters $\mu_{c}$ and $\Sigma_{c}$ describe a PSA model of type $\gamma_{c}$. Therefore, the number of free parameters is $\kappa(\gamma_{c})$, as defined in Equation (2). Since neither the means nor the covariance matrices are shared between the mixture components, the number of parameters for $\mu_{1},\Sigma_{1},\dots,\mu_{C},\Sigma_{C}$ is simply the sum $\sum_{c=1}^{C}\kappa(\gamma_{c})$. Eventually, since the mixture proportions are independent from the means and covariance matrices, the total number of parameters is $\kappa(\gamma)=C-1+\sum_{c=1}^{C}\kappa(\gamma_{c})$}. ∎
@@ A.2 Theorem 2: E-step of the EM
Let us define the cost function $K_{c}$ similarly to [bouveyron_high-dimensional_2007] via the following relationship: $\ln(\pi_{c}\,\mathcal{N}(\,x\mid\mu_{c},\Sigma_{c}\,))=-\frac{K_{c}(x)+p\ln 2\pi}{2}$. Re-injecting $K_{c}$ in the posterior probabilities yields
@@ A.3 Theorem 3: M-step of the EM
The solutions to such problems are given in the maximum likelihood estimation theorem of [szwagier_curse_2025]. One has $\hat{\mu}_{c}=\frac{1}{n\hat{\pi}_{c}}\sum_{i=1}^{n}t_{ci}\,x_{i}$ and $\hat{\Sigma}_{c}=\sum_{k=1}^{d_{c}}\hat{\lambda}_{ck}(t_{c},\gamma_{c})\,\hat{\Pi}_{ck}(t_{c},\gamma_{c})$, where $\hat{\lambda}_{ck}$ and $\hat{\Pi}_{ck}$ are the block-averaged eigenvalues and eigenspaces of $S_{c}$. ∎
@@ A.4 Theorem 4: monotonicity of the CPEM
The two previous equations (45, 46) simply mean that a classical EM iteration for the MPSA model of types $\gamma$ increases the expected complete-data penalized log-likelihood $\mathbb{E}_{\,y\mid x,\theta\,}[\ln\mathcal{L}_{\mathrm{c}}(\cdot)]-\alpha\,\kappa(\gamma)$, and that an adapted change of types from $\gamma$ to $\gamma^{\prime}$ increases this quantity even more. Using Gibbs’ inequality as classically done in the proof of the EM monotonicity [dempster_maximum_1977], one gets
@@ A.5 Proposition 5: GMM-based single-image denoising
[…]
@@ Appendix B Experiments
B.1 Penalty factor $\alpha$
In all the experiments, we set the penalty factor to $\alpha=\ln(n)/2$. This corresponds to a BIC-like penalized log-likelihood (23), which therefore enjoys theoretical foundations and asymptotic consistency guarantees [schwarz_estimating_1978, keribin_consistent_2000, bai_consistency_2018]. Notably, as we can see it in the density estimation experiments of Section 5.4, the estimated mixture parameters converge to the true ones when the number of samples goes to infinity. We also chose this criterion because this is a very classical one in model-based clustering (see, for instance, scrucca_mclust_2016, bouveyron_model-based_2019).
Alternatives criteria could be the Akaike information criterion (AIC) [akaike_new_1974] or the integrated completed likelihood (ICL) [biernacki_assessing_2000]. These criteria might not enjoy the same theoretical guarantees and result in different empirical results. Notably, the AIC ($\alpha=1$) is known to favor overparameterized models with respect to the BIC ($\alpha=\ln(n)/2$). While AIC and BIC directly express under the form of Equation (23)—i.e., as a log-likelihood penalized by the number of samples—ICL has a slightly different expression, which makes it non-integratable in the current framework, especially if we want monotonicity guarantees such as the ones of Theorem 4.
To account for the influence of the penalty factor, we report in Figure 9 the evolution of the number of parameters of the learned mixture model (via the CPEM algorithm) as a function of $\alpha$. [^12] The experimental details are the following ones: the data is generated from an MPSA density with $n=1000$, $p=20$, $C=3$, $(\gamma_{1},\gamma_{2},\gamma_{3})=((5,15),(5,5,10),(5,5,5,5))$}, $(\pi_{1},\pi_{2},\pi_{3})=(0.4,0.3,0.3)$, $\mu_{c}=10$, $(\lambda_{11},\lambda_{21},\lambda_{31})=(3,2,1)$, and $\lambda_{cp}/\lambda_{c1}=0.1\,\forall c$. The number of parameters is therefore $\kappa(\gamma)=421$. The experiments are repeated independently 10 times.
Figure 9: Influence of the regularization hyperparameter $\alpha$ on the number of parameters of the inferred mixture models. The true number of parameters is $\kappa(\gamma)=421$.
We see that the number of parameters indeed decreases with the penalty factor $\alpha$, going from the highest complexity (corresponding to the full GMM) to the lowest one (corresponding to the spherical GMM). The true number of parameters lies in between the ones selected by the BIC (underparameterized) and the AIC (overparameterized).
While the AIC and the BIC do not depend on the dimension $p$, alternative criteria such as the AICc [hurvich_regression_1989]—a small-sample version of the AIC—do. Some additional empirical results about the impact of the model selection criterion on the number of parameters of the PSA models are given in the appendix of [szwagier_curse_2025]. The AICc tends to select even more parsimonious models than the BIC for small-to-moderate sample sizes.
B.2 Number of components $C$
In contrast to many works from the GMM literature that reduce the number of mixture parameters by reducing the number of components $C$, our work controls parsimony through the covariance structure of the mixture components. While $C$ has a linear influence on the number of parameters, the eigenvalue multiplicities $\gamma_{ck}$ have a quadratic influence on the number of parameters (see Equation (2)). Therefore, in high dimensions, we believe that it is much more interesting to reduce the number of parameters via the eigenvalue multiplicities than via the number of components (unless $C$ is very large).
Except for the image denoising experiments in Section 6, this paper voluntarily involves experiments where the number of components $C$ is well specified—i.e., equal to the true number of clusters. This choice is motivated by the willingness to highlight the impact of the covariance structure on the learned mixtures—that is the main message of the paper—rather than the impact of the number of clusters.
In the current version of the code, the number of components $C$ can be updated during learning via several simple methodologies, for instance removing empty clusters, or the ones with less than a prespecified number of points. $C$ can also be estimated a posteriori, as commonly done, via model selection criteria such as the BIC (see, for instance, bouveyron_model-based_2019). As an illustration, we compare in Figure 10 the penalized log-likelihood of an MPSA-U model with $C=2$, $C=3$ and $C=4$, under the setting of the previous experiment—with three clusters.
Figure 10: Influence of the number of components $C$ on the penalized log-likelihood, under the MPSA-U model. The true number of clusters is $C=3$.
As we can see, the underfitted model ($C=2$) always has a significantly smaller penalized log-likelihood than the other models. Surprisingly, the overfitted model ($C=4$) achieves a significantly larger penalized log-likelihood than the true one ($C=3$) for $n\lessapprox 100$. Looking more closely at the estimated parameters, it seems that under the overfitted model, two clusters are well estimated, while the third one is split in two. This overparameterization naturally yields a larger log-likelihood, which is apparently not compensated by the penalty on the number of parameters—notably since the two overfitting components have lower number of parameters than the true one. Therefore, penalized log-likelihood criteria such as the BIC may have some limitations in selecting the true number of components, when the number of samples is small. For $n\gtrapprox 100$, the true model ($C=3$) gets a significantly larger penalized log-likelihood and is therefore selected over the overfitted one ($C=4$), testifying of the BIC’s asymptotic consistency.
Beyond the two simple ideas proposed above (removing unpopulated clusters and using model selection criteria), we believe that the choice of the number of clusters could be thoroughly addressed via Dirichlet-type priors on the mixture proportions. Such priors are commonly used in the context of overfitted mixture models (see, notably, the classical works of figueiredo_unsupervised_2002, rousseau_asymptotic_2011, malsiner-walli_model-based_2016). They enable to automatically remove underpopulated clusters within the M-step, without losing the theoretical monotonicity guarantees of the EM algorithm.
B.3 Patch size $s$ and number of components $C$ in denoising
Except from Table 3, the main text’s denoising experiments do not assess the impact of hyperparameters that are unrelated to the covariance structure on the results. This is the goal of this subsection.
The size of the patches naturally influences the denoising results. Intuitively, if the patches are too small, then they all look similar, while if they are too large, then they all look different—especially in textured zones. This has notable repercussions on the denoising performances, since the investigated methods precisely rely on a faithful modeling of the groups of similar patches extracted from one image (see Section 6.1). Figure 11 reports the influence of the patch size on the quality of denoising (PSNR). We can see that it first increases with the patch size, up until approximately $8\times 8$ (which is the size chosen in the experiments of Section 6) before decreasing. We can also see that the computing time naturally increases with the patch size, since the dimension of the dataset ($s^{2}$) increases quadratically with the patch size $s$. Therefore, beyond the denoising PSNR, the running time is another metric to take into account for choosing the patch size.
Similarly, the number of components $C$ does influence the denoising results. Intuitively, the goodness-of-fit increases with the number of components $C$. If the latter is too small, then we are in a situation of underfitting; the noise is overestimated and the denoising leads to a blurry image. In contrast, if the number of components is too large, then we are in a situation of overfitting; the noise is underestimated and the denoising does not modify much the input patch, resulting in a sharp yet noisy image; the limit case being for $C=n$, where each point is in its own cluster, the estimated noise is zero and the denoising leaves the noisy patches unchanged. These intuitions are (partially) confirmed empirically in Figure 11, which reports the influence of the number of clusters on the PSNR and running times. [^13] The decrease in denoising PSNR for large $C$ has been observed in houdard_high-dimensional_2018, but it appears to require an extremely large number of components to be significant, therefore we chose not to report it in Figure 11. Note that the running time increases linearly with $C$, therefore it might impact the practical choice of $C$ beyond the denoising performance. Notably, MPSA-U already reaches satisfying PSNRs for $C=5$.
Figure 11: Influence of the patch size $s$ (top) and the number of components $C$ (bottom) on the denoising quality (PSNR) and running time, under the MPSA-U model.
B.4 Truncated SVD
The code for MPSA inference has been partially optimized in order to run in practical time. However, it could largely enjoy more engineering, especially for very large dimensions. One potential improvement is the truncated SVD (TSVD) evoked in Section 3.3.
The current M-step for MPSA inference implies a full EVD of the responsibility-weighted covariance matrices $S_{c}$. Indeed, all the eigenvalues are needed to select the appropriate eigenvalue multiplicities. However, not all the eigenvectors are needed, but just the leading ones (as explained in Section 3.3). Hence, an idea can be to first compute all the eigenvalues (without the associated eigenvectors), then choose the appropriate eigenvalue multiplicities $\hat{\gamma}_{c}=(\hat{\gamma}_{c1},\dots,\hat{\gamma}_{cd_{c}})$ and finally compute the $\hat{q}_{c}=p-\hat{\gamma}_{cd_{c}}$ leading eigenvectors (associated to the $\hat{q}_{c}$ largest eigenvalues) via truncated SVD. While such a trick should not change the outcome in theory—with respect to a full EVD—it actually does in practice, since the truncated SVD relies on different solvers from the classical EVD. More details can be found in the documentation of the scipy package [virtanen_scipy_2020]: see scipy.linalg.eigh for the eigenvalue decomposition and scipy.sparse.linalg.svds for the truncated SVD. The randomized SVD of halko_finding_2011 is another popular solver.
Another simple idea is to impose an upper bound on the intrinsic dimensions of the mixture components (i.e. $\hat{q}_{c}\leq q_{\mathrm{max}}\,\forall c$). In that case, one can directly run a truncated SVD with $q_{\mathrm{max}}$ components without the need to compute all the eigenvalues a priori. This approach can be expected to be more efficient than the previous one, as it does not require to compute all the eigenvalues first. However, imposing such an upper bound on the intrinsic dimensions restricts the set of candidate models and may therefore lead to suboptimal results.
To compare these different approaches, we sample $n=1000$ points from a multivariate Gaussian distribution with zero mean and covariance $\mathrm{diag}\left(10,8,6,4,2,0.1^{p-5}\right)$, for varying $p$. We display the running times of the different approaches in Figure 12.
Figure 12: Comparison of the running times for different eigenvalue decomposition methodologies. EVD: full eigenvalue decomposition of the covariance matrix. SVD: full singular value decomposition of the data matrix. TSVD: truncated SVD of the data matrix, with five components (random: randomized SVD from [halko_finding_2011]; arpack: scipy solver which uses an Arnoldi iteration). EVD + TSVD: eigenvalue decomposition of the covariance matrix (without the eigenvectors) followed by TSVD with five components.
We see that the full SVD on the data matrix is overall slower than the other methodologies, including the full EVD on the covariance matrix. For very large dimensions ($p\gtrapprox 1000$), the methods involving first an eigenvalue decomposition without the eigenvectors and second a truncated SVD become significantly faster than the full EVD. The methods directly imposing an upper bound on the intrinsic dimensions are, for their part, significantly faster than the previous ones. Hence, we recommend integrating TSVD-based approaches in very large dimensions and directly imposing an upper bound on the intrinsic dimensions if speed and parsimony matter more than precise inference.
B.5 Benchmark
In our benchmark experiments, we mostly focus on comparing our MPSA models to the classical full and spherical GMMs—and to the less-classical HDDC [bouveyron_high-dimensional_2007] and HDMI [houdard_high-dimensional_2018]. This seemingly restrictive choice is actually well motivated: we want to assess the impact of equalizing close eigenvalues in covariance matrices on mixture modeling. We do not aim to achieve state-of-the-art results on density estimation, clustering and denoising tasks. Therefore, we believe that the best way to convey the main message of the paper (that is the interest of equalizing close eigenvalues in high-dimensional, low-sample settings) is to compare our models to baseline models that are particular cases of ours, and whose distinctive feature is the covariance structure alone. Indeed, if we compared our models to other baseline models (for density estimation, clustering or denoising), then the difference in performance could be explained by other features than the covariance structure, which would drown the core message of the paper.
B.6 Additional details
@@
This subsection provides more details about the experiments. The code is written in Python 3.9 and run on an Intel® Core™ i7-10850H CPU with 16 GB of RAM. The code for the EM algorithm is a modification of Mathieu Fauvel’s code (license: GPL-3.0), itself inspired from the R package of [berge_hdclassif_2012] (license: GPL-2). Some parts of the code related to PSA inference and model selection were inspired from the code of [szwagier_curse_2025] (license: MIT). The code for the denoising pipeline is a modification of Antoine Houdard’s code (license: MIT).
@@
The used packages are the following: scikit-learn [pedregosa_scikit-learn_2011] (license: BSD 3), numpy [harris_array_2020] (license: BSD), scipy [virtanen_scipy_2020] (license: BSD 3), matplotlib [hunter_matplotlib_2007] (license: custom), ucimlrepo (license: MIT), scikit-image [walt_scikit-image_2014] (license: BSD 3). The code for sampling from a skew-normal distribution comes from Gregory Gundersen’s blog post (license: not specified).
@@
Some datasets come from the UCI Machine Learning repository: wine [aeberhard_wine_1991] (license: CC BY 4.0), breast [wolberg_breast_1995] (license: CC BY 4.0) and ionosphere [sigillito_ionosphere_1989] (license: CC BY 4.0). The Simpson image, which is notably used in [houdard_high-dimensional_2018] comes from the website page of Antoine Houdard (license: not specified).